# FlowBot++: Learning Generalized Articulated Objects Manipulation via Articulation Projection

### v3

@@ 2 Related Work
Articulated Object Manipulation: Manipulation of unseen articulated objects and other objects with non-rigid properties remains an open research area due to the objects’ complex geometries and kinematics. Previous work proposed manipulating such objects using analytical methods, such as the immobilization of a chain of hinged objects, constraint-aware planning, and relational representations [7, 3, 8]. With the development of larger-scale datasets of articulated objects [9, 10], several works have proposed learning methods based on large-scale simulation, supervised visual learning and visual affordance learning [11, 12, 13]. Several works have focused on visual recognition and estimation of articulation parameters, learning to predict the pose [4, 5, 14, 15, 16] and identify articulation parameters [1, 17] to obtain action trajectories. Statistical motion planning methods could also be applied in the scenarios of complex hinged objects manipulation [18, 19, 20]. More closely related to our work is that of Eisner et al. [6], which learns to predict per-point motion of articulated objects under instantaneous motion. While the per-point motion affordance learning of Eisner et al. [6] outperforms previous approaches, it does not explicitly model the articulation parameters. Thus, the affordance needs to be estimated in each time step, yielding jerky robot trajectories. Concurrently with our work, Nie et al. [2] proposed a similar idea of estimating articulation parameters from interactions to guide downstream manipulation tasks. However, our method does not require a priori interactions with the manipulated objects.
Flow for Policy Learning: Optical flow [21] is used to estimate per-pixel correspondences between two images for object tracking and motion prediction and estimation. Current state-of-the-art methods for optical flow estimation leverage convolutional neural networks [22, 23, 24]Dong et al. [25], Amiranashvili et al. [26] use optical flow as an input representation to capture object motion for downstream manipulation tasks. Weng et al. [27] use flow to learn a policy for fabric manipulation. Previous work generalized optical flow beyond pixel space to 3D for robotic manipulation [6, 28, 29]. We learn a 3D representation for articulated objects, Articulation Flow, which describes per-point correspondence between two point clouds of the same object and Articulation Projection, a dense projection representation of the object’s articulation parameters.
@@ 4.1 A New Representation of Articulated Objects
where $\boldsymbol{\omega}$ is a unit vector that represents the axis direction and $v$ represents the origin of the axis and $\mathbf{I}$ represents a 3x3 identity matrix. Mathematically, $r_{p}$ is calculated as the difference between the vector $\overrightarrow{vp}$ and its projection onto $\vec{v}$, which is equivalently a perpendicular vector from each point to the axis of rotation defined by direction $\boldsymbol{\omega}$ and origin $v$, which is represented using the blue vectors in Fig. 2. As we will see, Articulation Projection can be used to predict a longer articulation trajectory while still inheriting the generalization properties of prior work [6].
@@ 4.3 Jointly Learning Articulation Flow and Articulation Projection
where $\oplus$ represents concatenation. We train FlowProjNet via supervised learning with this loss. We describe the method to obtain ground-truth labels in Appendix E. We implement FlowProjNet using a segmentation-style PointNet++ [30] as a backbone, and we train a single model across all categories, using a dataset of synthetically-generated data tuples based on the ground-truth kinematics provided by the PartNet-Mobility dataset [10]. Details of our dataset construction and model architecture can be found in Appendix E.
@@ 5.2 Real-World Experiments
We conduct qualitative real-world experiments to assess the sim2real capabilities of FlowBot++. We run FlowBot++ trained in simulation directly on denoised point clouds collected on real-world objects [32] without any retraining. We test our method on 6 real-world prismatic (Drawer) and revolute (Fridge, Oven, Toilet, Microwave, Trashcan) objects. In real-world trials, we provide hand-labeled segmentation masks to the network, which are used to filter and aggregate the results. We use a heuristically-defined grasping policy to select a grasp point and execute the planned trajectories on a Sawyer Black Robot equipped with a parallel-jaw gripper. Following the robot controller design in [28], once a contact point is chosen, we could transform the contact point using Eq. 5 because the gripper is now modeled as rigidly attached to the contact point, forming a full robot end-effector trajectory. Example predictions are shown in Fig. 6, in which all points on the segmented part are collectively transformed using Eq. 5, showing a full trajectory of the articulation. Real-world experiments show promising results that our networks transfer the learned 3D geometry information to real-world data. We show side-by-side comparisons between FlowBot++ and FlowBot3D on our website. FlowBot3D [6] completely failed on real-world oven, while FlowBot++ is able to predict a feasible full-trajectory to open the oven door with only one planning step. We also observe that FlowBot++ is able to produce smoother trajectories without undesired motions and movements of the objects and finish the execution in a shorter time frame, corroborating our findings in simulation. Details of the real-world experiments and system implementation are documented in Appendix F.
@@ Appendix C Baselines
  • Normal Direction: We use off-the-shelf normal estimation to estimate the surface normals of the point cloud using Open3D [33]. To break symmetry, we align the normal direction vectors to the camera. At execution time, we first choose the ground-truth maximum-flow point and then follow the direction of the estimated normal vector of the surface.
@@ E.1 Datasets
To evaluate our method in simulation, we implement a suction gripper in the PyBullet environment, which serves as a simulation interface for interacting with the PartNet-Mobility dataset [10]. The PartNet-Mobility dataset contains 46 categories of articulated objects; following UMPNet [12], we consider a subset of PartNet-Mobility containing 21 classes, split into 11 training categories (499 training objects, 128 testing objects) and 10 entirely unseen object categories (238 unseen objects). Several objects in the original dataset contain invalid meshes, which we exclude from evaluation. We train our models (FlowProjNet and baselines) exclusively on the training instances of the training object categories, and evaluate by rolling out the corresponding policies for every object in the dataset. Each object starts in the “closed” state (one end of its range of motion), and the goal is to actuate the joint to its “open” state (the other end of its range of motion). For experiments in simulation, we include in the observation $O_{t}$ a binary part mask indicating which points belong to the child joint of interest.

### v4

@@ 2 Related Work
Articulated Object Manipulation: Manipulation of unseen articulated objects and other objects with non-rigid properties remains an open research area due to the objects’ complex geometries and kinematics. Previous work proposed manipulating such objects using analytical methods, such as the immobilization of a chain of hinged objects, constraint-aware planning, and relational representations [7, 3, 8, 9]. With the development of larger-scale datasets of articulated objects [10, 11, 12], several works have proposed learning methods based on large-scale simulation, supervised visual learning and visual affordance learning [13, 14, 15, 16]. Several works have focused on visual recognition and estimation of articulation parameters, learning to predict the pose [4, 5, 17, 18, 19, 20] and identify articulation parameters [1, 21] to obtain action trajectories. Statistical motion planning methods could also be applied in the scenarios of complex hinged objects manipulation [22, 23, 24]. More closely related to our work is that of Eisner et al. [6], which learns to predict per-point motion of articulated objects under instantaneous motion. While the per-point motion affordance learning of Eisner et al. [6] outperforms previous approaches, it does not explicitly model the articulation parameters. Thus, the affordance needs to be estimated in each time step, yielding jerky robot trajectories. Concurrently with our work, Nie et al. [2] proposed a similar idea of estimating articulation parameters from interactions to guide downstream manipulation tasks. However, our method does not require a priori interactions with the manipulated objects.
Flow for Policy Learning: Optical flow [25, 26, 27] is used to estimate per-pixel correspondences between two images for object tracking and motion prediction and estimation. Current state-of-the-art methods for optical flow estimation leverage convolutional neural networks [28, 29, 30]Dong et al. [31], Amiranashvili et al. [32] use optical flow as an input representation to capture object motion for downstream manipulation tasks. Weng et al. [33] use flow to learn a policy for fabric manipulation. Previous work generalized optical flow beyond pixel space to 3D for robotic manipulation [6, 34, 35, 36]. We learn a 3D representation for articulated objects, Articulation Flow, which describes per-point correspondence between two point clouds of the same object and Articulation Projection, a dense projection representation of the object’s articulation parameters.
@@ 4.1 A New Representation of Articulated Objects
where $\boldsymbol{\omega}$ is a unit vector that represents the axis direction and $v$ represents the origin of the axis and $\mathbf{I}$ represents a 3x3 identity matrix. Mathematically, $r_{p}$ is calculated as the difference between the vector $\overrightarrow{vp}$ and its projection onto $\vec{v}$, which is equivalently a perpendicular vector from each point to the axis of rotation defined by direction $\boldsymbol{\omega}$ and origin $v$, which is represented using the blue vectors in Fig. 2. As we will see, Articulation Projection can be used to predict a longer articulation trajectory while still inheriting the generalization properties of prior work [6, 37, 38].
@@ 4.3 Jointly Learning Articulation Flow and Articulation Projection
where $\oplus$ represents concatenation. We train FlowProjNet via supervised learning with this loss. We describe the method to obtain ground-truth labels in Appendix E. We implement FlowProjNet using a segmentation-style PointNet++ [39, 40] as a backbone, and we train a single model across all categories, using a dataset of synthetically-generated data tuples based on the ground-truth kinematics provided by the PartNet-Mobility dataset [11, 41]. Details of our dataset construction and model architecture can be found in Appendix E.
@@ 5.2 Real-World Experiments
We conduct qualitative real-world experiments to assess the sim2real capabilities of FlowBot++. We run FlowBot++ trained in simulation directly on denoised point clouds collected on real-world objects [20, 43] without any retraining. We test our method on 6 real-world prismatic (Drawer) and revolute (Fridge, Oven, Toilet, Microwave, Trashcan) objects. In real-world trials, we provide hand-labeled segmentation masks to the network, which are used to filter and aggregate the results. We use a heuristically-defined grasping policy to select a grasp point and execute the planned trajectories on a Sawyer Black Robot equipped with a parallel-jaw gripper. Following the robot controller design in [34], once a contact point is chosen, we could transform the contact point using Eq. 5 because the gripper is now modeled as rigidly attached to the contact point, forming a full robot end-effector trajectory. Example predictions are shown in Fig. 6, in which all points on the segmented part are collectively transformed using Eq. 5, showing a full trajectory of the articulation. Real-world experiments show promising results that our networks transfer the learned 3D geometry information to real-world data. We show side-by-side comparisons between FlowBot++ and FlowBot3D on our website. FlowBot3D [6] completely failed on real-world oven, while FlowBot++ is able to predict a feasible full-trajectory to open the oven door with only one planning step. We also observe that FlowBot++ is able to produce smoother trajectories without undesired motions and movements of the objects and finish the execution in a shorter time frame, corroborating our findings in simulation. Details of the real-world experiments and system implementation are documented in Appendix F.
@@ Appendix C Baselines
  • Normal Direction: We use off-the-shelf normal estimation to estimate the surface normals of the point cloud using Open3D [44, 45]. To break symmetry, we align the normal direction vectors to the camera. At execution time, we first choose the ground-truth maximum-flow point and then follow the direction of the estimated normal vector of the surface.
@@ E.1 Datasets
To evaluate our method in simulation, we implement a suction gripper in the PyBullet environment, which serves as a simulation interface for interacting with the PartNet-Mobility dataset [11, 46]. The PartNet-Mobility dataset contains 46 categories of articulated objects; following UMPNet [14], we consider a subset of PartNet-Mobility containing 21 classes, split into 11 training categories (499 training objects, 128 testing objects) and 10 entirely unseen object categories (238 unseen objects). Several objects in the original dataset contain invalid meshes, which we exclude from evaluation. We train our models (FlowProjNet and baselines) exclusively on the training instances of the training object categories, and evaluate by rolling out the corresponding policies for every object in the dataset. Each object starts in the “closed” state (one end of its range of motion), and the goal is to actuate the joint to its “open” state (the other end of its range of motion). For experiments in simulation, we include in the observation $O_{t}$ a binary part mask indicating which points belong to the child joint of interest.